# Mitigate the Gap: Investigating Approaches for Improving Cross-Modal Alignment in CLIP

### v2

@@ Abstract
Contrastive Language–Image Pre-training (CLIP) has manifested remarkable improvements in zero-shot classification and cross-modal vision-language tasks. Yet, from a geometrical point of view, the CLIP embedding space has been found to have a pronounced modality gap. This gap renders the embedding space overly sparse and disconnected, with different modalities being densely distributed in distinct subregions of the hypersphere. In this work, we aim at answering two main questions: 1. Does sharing the parameter space between the multi-modal encoders reduce the modality gap? 2. Can the gap be mitigated by pushing apart the uni-modal embeddings via intra-modality separation? We design AlignCLIP, in order to answer these questions and show that answers to both questions are positive. Through extensive experiments, we show that AlignCLIP achieves noticeable enhancements in the cross-modal alignment of the embeddings, and thereby, reduces the modality gap, while maintaining the performance across several downstream evaluations, such as zero-shot image classification, zero-shot multi-modal retrieval and zero-shot semantic text similarity. The source code for reproducing our experiments is available at https://github.com/sarahESL/AlignCLIP.
@@ 1 Introduction
Through extensive experiments, Liang et al. (2022) and Tyshchuk et al. (2023) showed that the modality gap is caused by a combination of the model initialization and the contrastive loss optimization. Recently, Schrodi et al. (2024) showed that the driving factor behind the modality gap is the information-imbalance between the two modalities, i.e., texts in the training datasets usually have less detailed information in comparison to their pairing images. Furthermore, Jiang et al. (2023) show that exact modality alignment in multi-modal embeddings is suboptimal for downstream tasks.
Attempting to improve cross-modal alignment and reducing the modality gap in CLIP without losing performance in downstream tasks, we propose AlignCLIP, in the quest for answering two main questions: 1. Does sharing the parameter space between the multi-modal encoders reduce the modality gap? 2. Can the gap be further mitigated by pushing apart the uni-modal embeddings via intra-modality separation?
We answer the first question by sharing the transformer encoder and the projection layer in the vision and language encoders, and observe that it already results in noticeable cross-modal alignment improvements. As for the second question, we introduce the Intra-Modality Separation objective that encourages the embeddings within each modality to be expanded and pushed towards the opposite modality, i.e., text embeddings get pushed apart, towards image embeddings, and at the same, image embeddings get pushed apart and moved towards text embeddings. As a result, each modality gets moved towards the embeddings from the opposite modality and thereby, the modality gap can get reduced. Our experimental results support that the two aforementioned refinements show substantial improvement of the cross-modal alignment while maintaining, and slightly improving, the zero-shot transfer performance across a wide variety of downstream tasks.
@@ 2 Background, Notations and Concepts
$\mathcal{L}_{v\rightarrow l}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]},\quad\mathcal{L}_{l\rightarrow v}=\frac{1}{N}\sum_{j=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,j}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]}{\sum_{i=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]},$ (1)
@@ 3.2 Intra-Modality Separation
As shown in Figure 1A, each modality resides in a distinct dense subregion of the CLIP embedding space. We hypothesize that this phenomenon is a direct result of the cross-modal contrastive objective in CLIP, which merely optimizes the relative distances of image embeddings and text embeddings. Instead, a more practical scenario is to also consider the relative in-modality distances, i.e., pairwise distances of text embeddings and pairwise distances of image embeddings. We refer to this property as Intra-Modality Separation (IMSep) and define an objective function that enforces reasonable separation within each modality.
IMSep is achieved by a vision to vision as well as a text to text contrastive loss, where image–text pairs are the positive samples and any pairwise combination of image–image and text–text are considered as negative samples, respectively:
$\mathcal{L}_{v\rightarrow v}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1,\atop{i\neq j}}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{v}}^{\,j})\mathbin{/}\tau]},\quad\mathcal{L}_{t\rightarrow t}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1,\atop{i\neq j}}^{N}\exp[(\vec{e}_{\mathrm{t}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]}.$ (6)
In Figure 3, our approach for obtaining IMSep is schematically summarized. Given a batch of encoded image–text pairs, IMSep creates $\hat{y}_{\mathrm{vsep}}$ and $\hat{y}_{\mathrm{tsep}}$ logits and then minimizes the cross-entropy loss over $Y$ and $\hat{y}_{\mathrm{vsep}}$ as well as $\hat{y}_{\mathrm{tsep}}$.
To this end, first the pairwise cosine similarities are calculated by:
$\mathcal{V}=E_{\mathrm{v}}E_{\mathrm{v}}^{\intercal},\quad\mathcal{T}=E_{\mathrm{t}}E_{\mathrm{t}}^{\intercal},$ (7)
where $\mathcal{V}\in\mathbb{R}^{b\times b}$ and $\mathcal{T}\in\mathbb{R}^{b\times b}$. One should notice that while enforcing intra-modality separation, some samples might indeed be semantically similar to each other and therefore, must not be separated immensely. To this end, we regularize each intra-modality separation with respect to the pairwise semantic similarity of the samples. In order to calculate the semantic similarity within the uni-modal samples, we invoke a pre-trained sentence encoder and encode each text as $\vec{e}_{\mathrm{s}}^{\,i}\in\mathbb{R}^{d}$. We denote the corresponding batch of semantically encoded texts as $E_{\mathrm{s}}\in\mathbb{R}^{b\times d}$, and proceed to calculate the pairwise semantic similarity $S\in\mathbb{R}^{b\times b}$ and the distance $D\in\mathbb{R}^{b\times b}$ of the texts:
@@
respectively. We rely on $\mathcal{D}$ in order to re-scale $\mathcal{V}$ and $\mathcal{T}$. The goal of this re-scaling is to enforce a smaller dot product of the encoded images and texts, respectively, if they are semantically similar, i.e., have a smaller distance in $D$. Conversely, the dot products in $\mathcal{V}$ and $\mathcal{T}$ obtain larger values when the text samples are semantically distant according to $D$. By this re-scaling, we seek to control the minimization of $(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{v}}^{\,j})$ as well as $(\vec{e}_{\mathrm{t}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})$ when the samples are semantically similar. This is because the re-scaling mechanism enforces reducing these in-modality dot product values, and therefore, the dot products do not get strongly minimized in the cross-entropy loss, since the values are already small. Re-scaling is performed by:
$\mathcal{V_{D}}=\mathcal{V}\odot\mathcal{D},\quad\mathcal{T_{D}}=\mathcal{T}\odot\mathcal{D}.$ (9)
where $\odot$ denotes the element-wise product, in which $\mathcal{V_{D}}\in\mathbb{R}^{b\times b}$ and $\mathcal{T_{D}}\in\mathbb{R}^{b\times b}$. Afterwards, we calculate $\mathcal{M}=E_{\mathrm{v}}E_{\mathrm{t}}^{\intercal}$ and mask the non-diagonal values by $\textrm{diag}(\mathcal{M})=\mathbb{I}\odot\mathcal{M}$. We then obtain the vision–vision logits $\hat{y}_{\mathrm{vsep}}\in\mathbb{R}^{b\times b}$ as well as text–text logits $\hat{y}_{\mathrm{tsep}}\in\mathbb{R}^{b\times b}$ by:
$\vspace*{-1mm}\hat{y}_{\mathrm{vsep}}=\exp(\tau)\cdot[\text{diag}(\mathcal{M})+\mathcal{V_{D}}],\quad\hat{y}_{\mathrm{tsep}}=\exp(\tau)\cdot[\text{diag}(\mathcal{M})+\mathcal{T_{D}}]$ (10)
Ultimately, we define the Intra-Modality Separation loss as:
$\mathcal{L}_{\textrm{IMsep}}=H(\hat{y}_{\mathrm{vsep}},Y)+H(\hat{y}_{\mathrm{tsep}},Y),$ (11)
@@ 4.1 Training Dataset and Setup
We used the Conceptual Caption 12M (CC12M) dataset, which has been similarly used in previous work (Goel et al., 2022; Mu et al., 2022; Li et al., 2022; Changpinyo et al., 2021), for pre-training the models. We set our setup quite similar to that of the original CLIP with the ViT-B-16 backend, in order to ensure a fair comparison of CLIP with SharedCLIP and AlignCLIP. The pre-trained semantic encoder utilized in AlignCLIP for re-scaling text–text and image–image cosine similarities is the SBERT all-mpnet-base-v2 model. In order to fairly compare the effectiveness of each model, we trained all of them from scratch using the CC12M dataset and the OpenCLIP implementation (Cherti et al., 2023; Ilharco et al., 2021). Each model was trained for 6 days using an NVIDIA H100 GPU with batch size 512 for 30 epochs using AdamW optimization with a starting learning rate of $1\times 10^{-3}$, cosine scheduler, 10,000 warmup steps, a weight decay of $0.1$, and an initial temperature value of 0.07. The output embedding dimensionality for both the vision and language modalities is set to 768. We used the checkpoint from the last epoch in our evaluations of downstream experiments. In AlignCLIP, we set $\alpha=1$ and $\beta=\frac{1}{2}$. In terms of ablation study, we also trained IMSep loss without sharing the parameters between vision and language encoders. In Section A.1, more details about the training setup is provided. Furthermore, in Section A.3, we repeat our experiment when using output embedding dimension of 512 and show the results.
@@ 4.2 Cross-Modal Alignment
We start by reporting and comparing the alignment scores when using CLIP, SharedCLIP, and AlignCLIP models on the validation sets from CC3M, MSCOCO as well as the ImageNet-1K, CIFAR-100, and CIFAR-10 datasets. Table 1 summarizes the corresponding alignment scores. For ablations, we also report the results when the IMSep loss in AlignCLIP is trained without sharing the parameters. We observe that the original CLIP model has relatively low alignment scores, varying within $[0.38,0.47]$, across all five datasets. These scores mean that the average angle between the paired image–text embeddings is a value between 61 and 68 degrees. In contrast, sharing the parameter space in SharedCLIP results in noticeable improvements of up to 0.17 in the alignment scores. As a result, the average angle between the paired image–text embeddings decreases to about 51 degrees. Furthermore, using AlignCLIP yields even better alignment scores, ranging from 0.58 to 0.67, and a decreased average angle of 47 degrees between the multi-modal paired samples. Furthermore, AlignCLIP without sharing parameters achieves the best alignment scores. These observations confirm that AlignCLIP improves the cross-modal alignment in CLIP and thereby, reduces the modality gap.
We also plot the distribution of cosine similarities of the positive samples from MSCOCO validation dataset when encoded using the original CLIP, SharedCLIP, and AlignCLIP in Figure 4. We find that using SharedCLIP noticeably shifts the distribution of the cosine similarity of positive samples towards higher similarity values. A higher similarity of positive samples, i.e., image–text pairs, means achieving greater cross-modal alignment, and thus a lower modality gap. Furthermore, AlignCLIP shifts the distribution even more to the right, resulting in higher similarities of positive samples, better cross-modal alignment, and a reduction of the modality gap.
Figure 4: Pairwise cosine similarities of positive samples in MSCOCO.
Model ImageNet-1K CIFAR-100 CIFAR-10 Food-101 Stanford Cars
@@
CLIP 31.4 58.7 28.1 55.9 61.5 95.6 35.2 63.3 11.6 36.5
SharedCLIP 32.1 59.7 26.4 54.7 56.9 95.2 32.8 61.2 10.7 35.4
AlignCLIP (no sharing) 28.7 55.8 23.1 49.7 50.1 91.4 33.1 61.8 8.4 30.7
AlignCLIP 32.4 60 31.2 61.8 66.4 96.9 33.1 63.3 11.8 37.1
@@
4.3 Zero-Shot Transfer

### v3

@@ Abstract
Contrastive Language–Image Pre-training (CLIP) has manifested remarkable improvements in zero-shot classification and cross-modal vision-language tasks. Yet, from a geometrical point of view, the CLIP embedding space has been found to have a pronounced modality gap. This gap renders the embedding space overly sparse and disconnected, with different modalities being densely distributed in distinct subregions of the hypersphere. In this work, we aim at answering three main questions: 1. Does sharing the parameter space between the multi-modal encoders reduce the modality gap? 2. Can the gap be mitigated by pushing apart the uni-modal embeddings via intra-modality separation? 3. How do these gap reduction approaches affect the downstream performance? We design AlignCLIP, in order to answer these questions and through extensive experiments, we show that AlignCLIP achieves noticeable enhancements in the cross-modal alignment of the embeddings, and thereby, reduces the modality gap, while improving the performance across several zero-shot and fine-tuning downstream evaluations. The source code for reproducing our experiments is available at https://github.com/sarahESL/AlignCLIP.
@@ 1 Introduction
Through extensive experiments, Liang et al. (2022) and Tyshchuk et al. (2023) showed that the modality gap is caused by a combination of the model initialization and the contrastive loss optimization. Recently, Schrodi et al. (2024) showed that the driving factor behind the modality gap is the information-imbalance between the two modalities, i.e., texts in the training datasets usually have less detailed information in comparison to their pairing images.
Attempting to reduce the modality gap, or equivalently, increase the cross-modal alignment in CLIP while enhancing the performance in downstream tasks, under the same amount of information-imbalance, we propose AlignCLIP. Ultimately, we answer three main questions: 1. With the same training dataset, i.e., under the same amount of information-imbalance between the two modalities, does sharing the parameter space between the multi-modal encoders reduce the modality gap? 2. Can the gap be further mitigated by pushing apart the uni-modal embeddings via intra-modality separation? 3. How do these gap reduction approaches affect the downstream performance?
We answer the first question by sharing the transformer encoder and the projection layer in the vision and language encoders, and observe that it already results in noticeable cross-modal alignment improvements as well as performance increase. As for the second question, we introduce the Intra-Modality Separation objective that encourages the embeddings within the visual modality to be expanded and pushed towards the language modality. As a result, each modality gets moved towards the embeddings from the opposite modality and thereby, the modality gap can get reduced. Our experimental results support that the two aforementioned refinements show substantial improvement of the cross-modal alignment while improving the zero-shot transfer performance across a wide variety of downstream tasks as well as linear probing for image classification and fine-tuning in cross-modal retrieval. As apposed to previous work that attempted at reducing the gap with naive isomorphic translations with respect to the distance of image-text pairs (Liang et al., 2022; Schrodi et al., 2024; Tyshchuk et al., 2023), which can hurt the distances of the unpaired samples, and therefore, distort the meaningful structure of the embedding space, AlignCLIP reduces the gap by modifications that improve the semantic structure of the latent modality space, as motivated by (Jiang et al., 2023).
@@ 2 Background, Notations and Concepts
$\mathcal{L}_{v\rightarrow l}=\frac{-1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]},\quad\mathcal{L}_{l\rightarrow v}=\frac{-1}{N}\sum_{j=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,j}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]}{\sum_{i=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]},$ (1)
@@ 3.2 Intra-Modality Separation
As shown in Figure 1A, each modality resides in a distinct dense subregion of the CLIP embedding space. We hypothesize that this phenomenon is a direct result of the cross-modal contrastive objective in CLIP, which merely optimizes the relative distances of image embeddings and text embeddings. The cross-modal contrastive loss alone is not sufficient for imposing meaningful distances within the uni-modal embeddings, i.e., pairwise distances of text embeddings and pairwise distances of image embeddings. Therefore, we define an objective function that enforces reasonable distances within each modality by separating the uni-modal embeddings that are semantically dissimilar. In other words, we impose a semantically-regularized Intra-Modality Separation (IMSep) in addition to the CLIP’s objective function.
IMSep is achieved by a vision to vision contrastive loss, where image–text pairs are the positive samples and any pairwise combination of image–image is considered as negative samples, respectively:
$\mathcal{L}_{v\rightarrow v}=\frac{-1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1,\atop{i\neq j}}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{v}}^{\,j})\mathbin{/}\tau]}.$ (6)
In practice, given a batch of encoded image–text pairs, IMSep creates the vision-vision $\hat{y}_{\mathrm{vsep}}$ logits and then minimizes the cross-entropy loss over $Y$ and $\hat{y}_{\mathrm{vsep}}$. To this end, first the pairwise cosine similarities of the images in the batch are calculated by:
$\mathcal{V}=E_{\mathrm{v}}E_{\mathrm{v}}^{\intercal},\quad\text{where}\quad\mathcal{V}\in\mathbb{R}^{b\times b}.$ (7)
One should notice that while enforcing intra-modality separation, by minimizing the denominator in Eq. 6, some samples might indeed be semantically similar to each other and therefore, must not be separated immensely. To this end, we regularize the intra-modality separation with respect to the pairwise semantic similarity of the samples. In order to calculate the semantic similarity within the image samples, we utilize their pairing texts as the semantic supervision signal, and invoke a pre-trained sentence encoder for encoding each text as $\vec{e}_{\mathrm{s}}^{\,i}\in\mathbb{R}^{d}$. We denote the corresponding batch of semantically encoded texts as $E_{\mathrm{s}}\in\mathbb{R}^{b\times d}$, and proceed to calculate the pairwise semantic similarity $S\in\mathbb{R}^{b\times b}$ and the distance $D\in\mathbb{R}^{b\times b}$ of the texts:
@@
respectively. We rely on $\mathcal{D}$ in order to re-scale $\mathcal{V}$. The goal of this re-scaling is to enforce a smaller dot product of the encoded images, if they are semantically similar, i.e., have a smaller distance in $D$. Conversely, the dot products in $\mathcal{V}$ obtain larger values when the text samples are semantically distant according to $D$. By this re-scaling, we seek to control the minimization of $(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{v}}^{\,j})$ when the samples are semantically similar. This is because the re-scaling mechanism enforces reducing these in-modality dot product values, and therefore, the dot products do not get strongly minimized in the cross-entropy loss, since the values are already small. Re-scaling is performed by:
$\mathcal{V_{D}}=\mathcal{V}\odot\mathcal{D},\quad\text{where}\quad\mathcal{V_{D}}\in\mathbb{R}^{b\times b}$ (9)
where $\odot$ denotes the element-wise product.
Afterwards, we calculate $\mathcal{M}=E_{\mathrm{v}}E_{\mathrm{t}}^{\intercal}$ and mask the non-diagonal values by $\textrm{diag}(\mathcal{M})=\mathbb{I}\odot\mathcal{M}$. We then obtain the vision–vision logits by:
$\vspace*{-1mm}\hat{y}_{\mathrm{vsep}}=\exp(\tau)\cdot[\text{diag}(\mathcal{M})+\mathcal{V_{D}}],\quad\hat{y}_{\mathrm{vsep}}\in\mathbb{R}^{b\times b}$ (10)
In Figure 3, our approach for obtaining IMSep is schematically summarized. Ultimately, we define the Intra-Modality Separation loss as:
$\mathcal{L}_{\textrm{IMsep}}=H(\hat{y}_{\mathrm{vsep}},Y),$ (11)
@@
Note that it is sufficient to define the Intra-Modality Separation function for only one of the modalities, image in our case, since the cross-modal objective defined by $\mathcal{L}_{\textrm{CRsep}}$ already behaves as a supervision for the other modality, text in our case, and enforces the Intra-Modality Separation in that modality as well. To better support this statement, we extensively experiment the effects of adding the similar Intra-Modality Separation on text embeddings in Section A.2 and show that it is sufficient to impose the Intra-Modality Separation on the image embeddings.
@@ 4.1 Training Dataset and Setup
We used the Conceptual Caption 12M (CC12M) dataset, which has been similarly used in previous work (Goel et al., 2022; Mu et al., 2022; Li et al., 2022; Changpinyo et al., 2021), for pre-training the models. We set our setup quite similar to that of the original CLIP with the ViT-B-16 backend, in order to ensure a fair comparison of CLIP with SharedCLIP and AlignCLIP. The pre-trained semantic encoder utilized in AlignCLIP for re-scaling image–image cosine similarities is the SBERT all-mpnet-base-v2 model. In order to fairly compare the effectiveness of each model, we trained all of them from scratch using the CC12M dataset and the OpenCLIP implementation (Cherti et al., 2023; Ilharco et al., 2021). Each model was trained for 6 days using an NVIDIA H100 GPU with batch size 512 for 30 epochs using AdamW optimization with a starting learning rate of $1\times 10^{-3}$, cosine scheduler, 10,000 warmup steps, a weight decay of $0.1$, and an initial temperature value of 0.07. The output embedding dimension for both the vision and language modalities is set to 768. We used the checkpoint from the last epoch in our evaluations of downstream experiments. In AlignCLIP, we set $\alpha=1$ and $\beta=\frac{1}{2}$. In Section A.1, more details about the training setup is provided.
Since our goal is to shed lights on a specific shortcoming of the original CLIP model, i.e., the modality gap problem, we compare our results to the original CLIP model in this paper. Further comparisons to other state-of-the-art models is not the focus of this study, as the goal is to investigate modifications that reduce the modality gap in CLIP without losing performance in downstream tasks, rather than achieving the state-of-the-art results.
@@
Figure 4: Cumulative distribution of pairwise cosine similarities of positive samples in MSCOCO.
@@ 4.2 Cross-Modal Alignment
We start by reporting and comparing the alignment scores when using CLIP, SharedCLIP, and AlignCLIP models on the validation sets from CC3M, MSCOCO as well as the ImageNet-1K, CIFAR-100, and CIFAR-10 test datasets. Table 1 summarizes the corresponding alignment scores. We observe that the original CLIP model has relatively low alignment scores, varying within $[0.38,0.47]$, across all five datasets. These scores mean that the average angle between the paired image–text embeddings is a value between 61 and 68 degrees. In contrast, sharing the parameter space in SharedCLIP results in noticeable improvements of up to 0.17 in the alignment scores. As a result, the average angle between the paired image–text embeddings decreases to about 51 degrees. Furthermore, using AlignCLIP yields even better alignment scores, ranging from 0.62 to 0.67, and a decreased average angle of 47 degrees between the multi-modal paired samples. These observations confirm that AlignCLIP improves the cross-modal alignment in CLIP and thereby, reduces the modality gap.
We also plot the cumulative distribution of cosine similarities of the positive samples from MSCOCO validation dataset when encoded using the original CLIP, SharedCLIP, and AlignCLIP in Figure 4. We find that using SharedCLIP noticeably shifts the distribution of the cosine similarity of positive samples towards higher similarity values. A higher similarity of positive samples, i.e., image–text pairs, means achieving greater cross-modal alignment, and thus a lower modality gap. Furthermore, AlignCLIP shifts the distribution even more to the right, resulting in higher similarities of positive samples, better cross-modal alignment, and a reduction of the modality gap.
Model ImageNet-1K CIFAR-100 CIFAR-10 Flowers-102 Stanford Cars
@@
CLIP 31.4 58.7 28.1 55.9 61.5 95.6 18 39.1 11.6 36.5
SharedCLIP 32.1 59.7 26.4 54.7 56.9 95.2 18.2 38.9 10.7 35.4
AlignCLIP 32.8 60.6 36.5 66.4 69.3 97.8 18.8 40.3 11.8 38.1
@@
Model ImageNet-1K CIFAR-100 CIFAR-10 Flowers-102 Stanford Cars
CLIP 50 62.6 85 71.5 42.2
SharedCLIP 51.2 63 85 74.4 40.5
AlignCLIP 51.5 67.4 87.2 76.8 45.6
4.3 Classification